OpenStudy (anonymous):

\[\int_0^2\int_0^3(x^2+y^2)\ dy\ dx\]

OpenStudy (anonymous):

24

OpenStudy (anonymous):

should be 26

OpenStudy (anonymous):

Yes, 26 after I worked it out

OpenStudy (calculator):

\[\int\limits _{-L}^L\int\limits _{-L}^L\left(x^2+y^2\right)dydx=\frac{8 L^4}{3}\]